# Defending against Reverse Preference Attacks is Difficult

### v1

@@ Abstract
While there has been progress towards aligning Large Language Models (LLMs) with human values and ensuring safe behaviour at inference time, safety-aligned LLMs are known to be vulnerable to training-time attacks such as supervised fine-tuning (SFT) on harmful datasets. In this paper, we ask if LLMs are vulnerable to adversarial reinforcement learning. Motivated by this goal, we propose Reverse Preference Attacks (RPA), a class of attacks to make LLMs learn harmful behavior using adversarial reward during reinforcement learning from human feedback (RLHF). RPAs expose a critical safety gap of safety-aligned LLMs in RL settings: they easily explore the harmful text generation policies to optimize adversarial reward. To protect against RPAs, we explore a host of mitigation strategies. Leveraging Constrained Markov-Decision Processes, we adapt a number of mechanisms to defend against harmful fine-tuning attacks into the RL setting. Our experiments show that “online" defenses that are based on the idea of minimizing the negative log likelihood of refusals—with the defender having control of the loss function—can effectively protect LLMs against RPAs. However, trying to defend model weights using “offline" defenses that operate under the assumption that the defender has no control over the loss function are less effective in the face of RPAs. These findings show that attacks done using RL can be used to successfully undo safety alignment in open-weight LLMs and use them for malicious purposes.
1 Introduction
It is well known that safety-aligned Large Language Models (LLMs) are vulnerable to harmful fine-tuning attacks (HFTAs) (Lermen et al., 2024; Yang et al., 2023; Zhan et al., 2024; Qi et al., 2024b). However, little is known about the robustness of safety guards during adversarial reinforcement learning from human feedback (RLHF) (Yi et al., 2024; Qi et al., 2024a) since its threat model and scope of attacks have not been fully formulated. In our study we develop the threat model of “reverse preference attacks” (RPAs) as an adversarial reinforcement learning setting i.e. high reward is given by an adversary to optimize for harmful policies. Figure 1 illustrates RPAs which involve an adversary flipping the preference label of an annotator such that a harmlessness preference dataset aligns with the goal of harmfulness such as willingness to give instructions on how to make a weapon. This is concerning, considering the popularity of RLHF as a training mechanism (Kaufmann et al., 2023) and the emergence of extensively used preference learning datasets (Bai et al., 2022; Ji et al., 2023; Dai et al., 2024; Köpf et al., 2023).
@@ 1 Introduction
Figure 1: Reverse preference attacks involve an adversary flipping the preference of an annotator.
Leveraging popular RLHF algorithms like DPO (Rafailov et al., 2023) and PPO (Schulman et al., 2017) for training-time attacks present a different threat model than supervised fine-tuning (SFT):
  1. 1.
    Harmlessness preference datasets may be subject to complete label flipping by an adversary. Data corruptions can also happen inadvertently, for example through human factors affecting data collectors or crowdworkers such as value misalignment, fatigue, or unclear understanding of labelling guidelines (Casper et al., 2023) (Table 2).
  2. 2.
    While reward models trained on such datasets would be subject to misspecification as a result of label noise, there is little work on how LLMs might optimize these or adversarial rewards generally . We find that effective defences either prevent exploration of high rewards or learn reward hacking which doesn’t result in harmful behaviour but breaks the models fluency (non-malicious reward hacking, Section 5.1).
  3. 3.
    RLHF mechanisms leverage policy learning constraints classically formulated as a KL-divergence between the logits of a policy model and a newly proposed policy. Qi et al. (2024a) regard this as a potential natural defence but we find that KL constraints are not generally indicators of defence success (Appendix F).
@@
In this paper, we propose Reverse Preference Attacks (RPA, Section 2), a class of RL-based attacks that exploit the above threat model. To investigate RPAs, we start with a vulnerability analysis that investigates how open-weight LLMs without additional defense mechanisms are vulnerable to RPA, based on varying ratios of preference flipping as well as complete preference flipping. We find that popular RLHF methods, like DPO and PPO, easily undo safety guards even in cases with as little as 25% of preference label flips. We then present a comprehensive empirical evaluation of defence mechanisms against RPA. Originally proposed in the context of HFTAs, we extend them through the framework of Constrained Markov Decision Processes (CMDPs) for Safe RL. We divide these methods into two categories. Online methods that apply some defence intervention during training provide an explicit constraint over the set of policies allowable during RLHF training. Offline methods apply some defence intervention before training such that the attack is made harder. We connect them with CMDPs by introducing the notion of implicit constraints, with the goal of removing harmful representations from the model weights themselves (Section 4.2).
@@
Experimentally, we show that only some methods of defence against harmful fine-tuning attacks are effective in the RPA setting. Generally, the only effective method across all methods is the online method introduced by Qi et al. (2024a) and the weight drift method introduced by Huang et al. (2024a). However, these methods suffer from requiring a very high penalty to the original loss function of the RLHF method which may impact its effectiveness of training on harmless tasks due to the underweighting the original loss function term compared to the defence loss term. Among offline defenses, Representation Noising (Rosati et al., 2024a, RepNoise) show promising results but fail to undo the effects of RPA satisfactorily. Given the proliferation of open-weight LLMs, our findings underscore the importance of using RPAs as a Red Teaming method for improving LLM safety, as well as investing in research on Blue Teaming methods in an offline setting where defenders can effectively and proactively harden an LLM before releasing its weights in the open.
2 Threat Model: Reverse Preference Attacks
Rosati et al. (2024b) introduced a threat model illustrating harmful fine-tuning attacks (HFTA) as a function of the attacker’s compute budget in terms of training steps taken. The attacker has unrestricted access and control of a safety-aligned model, and their goal is to use samples from a harmful dataset of prompts and responses to perform SFT so that the fine-tuned model is made capable of generating harmful responses.
@@ 2 Threat Model: Reverse Preference Attacks
Instead of SFT, we focus on a preference optimization setting. We start with a dataset $\mathcal{D}=\{x^{i},y^{i}_{w},y^{i}_{l}\}^{N}_{i=1}$ composed of potentially harmful prompts $x$ (such as How to make a molotov cocktail), as well as preferred ($y_{w}$) and dispreferred ($y_{l}$) samples—$w$ denoting win and $l$ denoting lose. The goal of RLHF is to use the preference $y_{w}\succ y_{l}$ to learn a reward model $R_{\phi}(y,x)$ over prompts and their responses. This reward model is used in a RL algorithm such as PPO (Schulman et al., 2017) to train a language model policy $\pi_{\theta}(\cdot|x)$ such that text generations are aligned with the reward model learned over preferences:
@@
Since we only will have a single-turn dialogue preference, this is modeled as a contextual bandit problem and there is no need to account for cumulative discounted reward over a sequence of actions.
A Reverse Preference Attack (RPA) is performed by reversing the preference data $y_{w}\prec y_{l}$—for example, by switching labels—inducing a "reverse" reward model $R_{\phi}^{\text{reverse}}(y,x)$. If the preference data is primarily over harmlessness, then this reward model can be used as an objective for learning harmful text generation policies by downstream RLHF algorithms.
Yi et al. (2024) introduced a special case of RPA in a restricted DPO setting.They showed “reverse DPO” was effective as HFTAs in undoing safety guards of LLMs. However, they neglected providing a full framework accounting for the unique threat of adversarial reward which we provide in this paper. In practice, an attacker would prefer to perform RPA over HFTA for the same reasons why a model developer might prefer to use RLHF over SFT: the desired behavioural policy (such as answering in a harmful manner) is easier to model as preference data or a reward signal than specific language samples for imitating with next-token prediction. To produce an RPA dataset, an attacker might leverage unique vulnerabilities such as attacking an online feedback collection system or exploiting a preference learning API.
3 Vulnerability Analysis
@@
To illustrate the effectiveness of RPAs, we assume a setting where the attacker has the capability to perform RLHF on an aligned language model using their own data—either by acquiring the weights of a model, or using a preference learning API exposed through a remote LLM provider. We perform our experiments on the safety-aligned open-weight model llama2-7b-chat (Touvron et al., 2023). As attack methods, we use HFTA (using SFT), and RPA (using DPO and PPO). As attack datasets, we use the harmlessness training sets of two widely used safety datasets: BeaverTails (Ji et al., 2023) and Safe RLHF (Dai et al., 2024). We flip a certain percentage of the harmfulness/preference labels of 1,000 randomly picked samples, and use each training algorithm to further train the original aligned LLM. A varying ratio of preference labels are flipped in order to simulate either attacker stealthiness or label noise.
@@
Table 1 presents experimental results for the case when 100% of the labels in an attack dataset are flipped. Specifically, it shows the mean probability of the harmfulness label of the answer to a held-out evaluation set of 100 harmful questions from each attack dataset using a harmfulness classifier. We observe that all three training algorithms are able to break safety alignment in the original model to a considerable degree on as little as 1,000 attack samples—RPA using DPO being the most effective. DPO and PPO use policy divergence constraints, $\beta=0.1$ and $kl=0.2$ respectively, with the original policy, i.e. the aligned model. We vary these constraints in Appendix F and find that policy divergence constraints with an aligned model are not an effective defence for DPO but may be effective for PPO which is likely because KL is used as a more direct constraint for PPO. Details of the experimental conditions of the attack, harmfulness measure and classifier, and datasets used are presented in Appendix B.
Attack No Attack HFTA–SFT RPA–DPO RPA–PPO
BeaverTails 0.06 $\pm$ 0.08 0.68 $\pm$ 0.38 0.70 $\pm$ 0.33 0.69 $\pm$ 0.30
Safe RLHF 0.07 $\pm$ 0.11 0.75 $\pm$ 0.29 0.83 $\pm$ 0.22 0.59 $\pm$ 0.33
@@
Next, we present the setting where labels are flipped at a given ratio $\rho\in[0,0.9]$ to simulate two realistic noise settings: (1) unintended data corruption due to label noise, and (2) adversarial data poisoning where an attacker hides a given number of labels. In the case of SFT, label flipping means we are using a harmful sample for causual language modeling instead of a harmless one. We use the same experimental setup as above, except that we also employ an unaligned model llama2-7b in order to understand the effects of vulnerability to HFTA and RPA during an alignment procedure. As shown in Table 2, in as little as 25% label flips for SafeRLHF we have successful attacks (0.51) which indicates that these models are vulnerable to label noise or stealthy attacks. RPA using DPO remains the most effective attack across noise levels and both datasets. Observe that DPO’s and PPO’s harmfulness does not always increase linearly with attack strength which due to the instability of training. PPO generally requires all labels to be flipped. As expected unaligned models are even more vulnerable than aligned ones which indicates that noisy harmlessness datasets could introduce significant issues during alignment training.
Model Dataset Attack 0.0 0.1 0.25 0.5 0.75 0.9
Aligned BeaverTails DPO 0.17 0.08 0.06 0.03 0.06 0.16
PPO 0.06 0.06 0.08 0.09 0.11 0.13
SFT 0.06 0.09 0.09 0.33 0.52 0.60
Safe RLHF DPO 0.13 0.28 0.51 0.76 0.82 0.83
PPO 0.07 0.06 0.17 0.12 0.38 0.13
SFT 0.37 0.41 0.51 0.60 0.65 0.69
Unaligned BeaverTails DPO 0.09 0.15 0.26 0.50 0.51 0.32
PPO 0.03 0.06 0.47 0.57 0.66 0.22
SFT 0.08 0.09 0.17 0.35 0.49 0.67
Safe RLHF DPO 0.12 0.26 0.55 0.76 0.80 0.84
PPO 0.06 0.09 0.40 0.59 0.44 0.59
SFT 0.38 0.36 0.45 0.54 0.62 0.69
@@
Table 2: Vulnerability of models under varying ratios (the columns) of harmless preference flips. The mean harmfulness scores are presented as above. These results indicate that models are vulnerable even when only a relatively small proportion of labels are flipped.
4 Defences
@@
To introduce defences against RPAs, we draw upon the Constrained Markov Decision Processes formalism from Dai et al. (2024, CMDP) and adapt a number of recently proposed defense techniques for HFTA (SFT) to the RLHF setting. In CMDP, the set of policies $\Pi_{\Theta}$ that are learnable in a typical policy learning algorithm, Equation 1, are restricted by a cost function $c$ and a cost threshold $b$ where the acceptable policy set is in $\Pi_{c}=\{\pi_{\theta}\in\Pi_{\Theta}|\>c(\pi_{\theta})\leq b)$. When the cost function is a harmfulness measure and the threshold is set to allow a degree of harm that is acceptable, this definition adapts the resistance condition for defences against HFTAs from Rosati et al. (2024b) to the RL setting. We connect this further to the attacker’s budget by using a measure of how long $\pi_{\theta}$ remains in $\Pi_{c}$ across training steps $t$ where the defender wants to prolong policy constraints under an attack.
@@
All of our defences adopt a variant of a constraint according to CMDP as summarized in Table 3. Broadly, we classify the defenses into two categories by defender capability assumptions: online and offline. Online defences assume that the defender does not have control of the training data but does have control over the training process. Online defences work by adopting different variants of explicit constraints on the training process. On the other hand, offline defences assume that the defender has applied before publicly releasing the weights of an LLM and can make no other interventions. Offline defences use implicit constraints that exist in the model weights themselves as demonstrated by their ability to prevent defended models from learning policies outside of the constraint set. Each defence is described in detail in Section B.1111Other viable defence settings such as data filtration are viable but outside of our scope of interest.
Defender Capability Constraint Defence Paper
Online Embeddings Vaccine (Huang et al., 2024b)
Weights Lisa (Huang et al., 2024a)
Rewards Safe RLHF (Dai et al., 2024)
Loss Security Vectors (Zhou et al., 2023)
Loss Refusal Loss (Qi et al., 2024a)
Offline Representation RepNoise (Rosati et al., 2024a)
Representation Circuit Breakers (Zou et al., 2024)
Representation RMU (Li et al., 2024b)
Meta-learned TAR (Tamirisa et al., 2024)
@@
Table 3: Taxonomy of defences candidates against RPAs showing that defences can be unified under the CDMP framework. The constraint column indicates how the harmlessness constraint is implemented.
4.1 Online Defences
We evaluate the online defences listed in Table 3 on the same attacks in Section 3. In addition to mean harmfulness scores, we use three additional metrics:
  1. 1.
    Helpfulness, measured by taking 100 samples of the preferred helpful questions from the Anthropic-HH helpfulness dataset (Bai et al., 2022) and using a helpfulness reward model (details in Appendix B) to rate generated responses to these questions. This measure is introduced to understand the degree to which the defence maintains general capabilities on harmless tasks.
  2. 2.
    Perplexity (PPL), measuring the fluency of the helpfulness answer above using gpt-2 (Radford et al., 2019).
  3. 3.
@@ 4.1 Online Defences
Table 4: Evaluation of online defences to HFTA and RPA variants. Missing metrics, indicated by dashes, indicate the training process could not complete due to sampling errors. Refusal loss and Lisa maintain the lowest harmfulness and highest helpfulness.
In Table 4 we find that across all methods Lisa and Refusal Loss are the most effective defences, maintaining low perplexity and high helpfulness. Security Vectors and Vaccine provide some defence against HFTA using SFT, but they do so at the cost of lower helpfulness scores. They also do not work in the other settings: Security Vectors is successfully attacked under RPA using both PPO and DPO, and Vaccine results in sampling errors during RPA/PPO due to numeric instability of the learned log probabilities or results in complete degeneration in the case of DPO, as evidenced by the method’s perplexity. For Safe RLHF, the method is limited to the PPO setting only as it is a reward shaping method and is not as effective (Safe RLHF dataset) or successfully attacked (BeaverTails). Interestingly, successful defences often have a very high KL divergence with the original policy. While KL divergence magnitudes are not necessarily informative of the actual distributional distances, since it is not a distance metric, this finding indicates that KL divergence with respect to the original policy model is not an effective indicator of preventing the undoing of safety alignment.
4.2 Offline Defences
Offline defences operate based on removing representations of harmfulness (see discussion in Section B.1). We ran a number of representation removal methods—as described in Table 3—as offline defences. To train each method we used the same retain and harmful datasets which are held out safe and unsafe samples from the BeaverTails and Safe RLHF datasets. We ran trainings for each method for 877 gradient steps using a batch size of $8$ for 7,016 samples (additional details available in Appendix B). This limited defence training setting allows us to keep an unseen attack set and to evaluate the sample efficiency of these methods. Note that unlike online defences, high perplexity and low helpfulness are not considered undesirable in offline defences, since we actually might want offline defences to result in broken models when subjected to harmful optimization pressure Henderson et al. (2023). Finally, we attempt three variants of the meta-learning defence TAR— TAR itself, combining the TAR and RepNoise losses during defence training (TAR + RepNoise), and applying TAR after RepNoise (RepNoise $\rightarrow$ TAR).

### v2

# Mitigating Unsafe Feedback with Learning Constraints

@@ Abstract
While there has been progress towards aligning Large Language Models (LLMs) with human values and ensuring safe behaviour at inference time, safety-guards can easily be removed when fine-tuned on unsafe and harmful datasets. While this setting has been treated extensively, another popular training paradigm, learning from unsafe feedback with reinforcement learning, has previously been unexplored. This is concerning due to the widespread deployment of feedback collection systems. We address this gap by providing an analysis of learning settings where feedback is adversarial and noisy, i.e. that unsafe samples are preferred over safe ones despite model developers goal to maintain safety. We find that safety-aligned LLMs easily explore unsafe action spaces through generating harmful text and optimize for adversarial reward indicating that current safety guards are not enough to prevent learning from unsafe feedback. In order to protect against this vulnerability, we adapt a number of both “implict” and “explicit” harmful fine-tuning defences to evaluate whether they are effective as learning constraints in an RL setting finding that no method is generally effective pointing to the need for more research in defences given the widespread adoption of methods designed to learn from feedback. We end the paper with the observation that some defences work by performing “harmless reward hacking” for which we provide a theoretical explanation drawn from the theory of Constrained Markov Decision Processes and provide some direction for future defence development.
Introduction
Safety guards of Large Language Models (LLMs) can easily be removed with fine-tuning on harmful datasets or even by accident (Lermen, Rogers-Smith, and Ladish 2024; Yang et al. 2023; Zhan et al. 2024; Qi et al. 2024b). Little is known about the robustness of safety guards during reinforcement learning from human feedback (RLHF) (see initial studies in Yi et al. 2024; Qi et al. 2024a). This is unfortunate since we are not sure whether safety guarded models might be vulnerable to learning from unsafe feedback which is made even more important by the widespread adoption of RLHF techniques and datasets including many AI applications which expose online feedback collection mechanisms, are built on crowdsourced data, or data from online repositories which are vulnerable to data poisoning (see Casper et al. (2023) for more on the real world significance of this threat).
In our study we focus on an adversarial reinforcement learning setting setting where unsafe and harmful text generation policies are rewarded. With the research question of: “Can we prevent LLMs from learning from unsafe feedback?” Our scope is exclusively focused on RLHF where the reward model is either implicit (e.g. Direct Policy Optimization (DPO) Rafailov et al. 2023) or explicitly trained from preference data that contains some number of unsafe samples that have been adversarially placed. We call this setting “Reverse Preference Attacks” (RPAs) to contrast with “Harmful fine-tuning attacks” (HFTAs) performed with supervised fine-tuning that have been discussed in the past (Rosati et al. 2024b).
@@ Introduction
Figure 1: Reverse preference attacks (RPAs) involve an adversary flipping the preference of an annotator.
To investigate RPAs, we start with a vulnerability analysis that investigates how LLMs without additional defence mechanisms are vulnerable to RPAs, based on varying ratios of unsafe samples in the feedback dataset. We find that popular RLHF methods, like DPO and PPO, can easily undo safety guards even in cases with as little as 25% unsafe preferences. We acknowledge this is too a large proportion of samples when stealthiness is important to the attacker as model developers would likely be able to detect and filter out these samples. However, we believe there is still cause for concern for example if systems are deployed in an online learning environment.
@@
We then present a comprehensive empirical evaluation of defence mechanisms against RPAs. Originally proposed in the context of HFTAs, we extend defences through the framework of Constrained Markov Decision Processes (CMDPs). We divide these methods into two categories. Online methods that apply some defence intervention during training and provide an explicit constraint over the set of policies allowable during RLHF training. Offline methods apply some defence intervention before training such that the attack is made harder or the safety guard is strengthened. We connect them with CMDPs by introducing the notion of implicit learning constraints.
@@
Experimentally, we show that only some methods attacks are able to mitigate RPAs and none are able to perfectly protect against learning from unsafe feedback. Generally, the only effective method across all methods is the online method introduced by Qi et al. (2024a) and the weight drift method introduced by Huang et al. (2024). However, these methods suffer from requiring a very high penalty to the original loss function of the RLHF method which may impact its effectiveness of training on harmless tasks due to the under weighting the original loss function term compared to the defence loss term. Among offline defences, most defences fails to protect against RPAs comprehensively. To mitigate this we provide a novel defence which combines Representation Noising (Rosati et al. 2024a, RepNoise) and (Tamirisa et al. 2024, TAR).
While, we do find that supervised fine-tuning is more effective and easier to perform attacks with, the significance of our results lie in the understanding of current LLMs vulnerabilities during feedback learning and whether they can be protected within that scope given that feedback collection has become an important part of ML systems. We end the paper with a reward analysis that shows a surprising result that defences which do mitigate optimizing a harmful reward model perform “harmless reward hacking”. That is defended models learn to get high reward while generating harmless, often dis-fluent, text sequences rather than harmful text sequences. We illustrate the utility of the CMDP framework by providing an explanation of why “harmless reward hacking” might be an expected result of trying to solve a CMDP in this setting. Future work may be able to leverage this insight for developing defences that result in ‘self-destructing models’ (Henderson et al. 2023) i.e. ones whose capabilities are destroyed under harmful optimization pressure.
Reverse Preference Attacks
@@
Our focus is a preference optimization setting. We start with a dataset $\mathcal{D}=\{x^{i},y^{i}_{w},y^{i}_{l}\}^{N}_{i=1}$ composed of potentially harmful prompts $x$ (such as How to make a molotov cocktail), as well as preferred ($y_{w}$) and dis-preferred ($y_{l}$) text samples— using standard notation (Dai et al. 2024) from RLHF we denote $w$ as preferred (called win since this sample wins over the lose sample) and $l$ as dis-preferred (lose). This is a pairwise rank order between text samples that is established by annotators. The goal of RLHF is to use the preference ordering expressed in the collected feedback to learn a reward model $R_{\phi}(y,x)$ over prompts and which is learned as a maximum likelihood estimator of the rank order of the preference dataset i.e. $\Pr(y_{w}\succ y_{l})$ This reward model is used in an RL algorithm such as PPO (Schulman et al. 2017) to train a language model policy $\pi_{\theta}$ such that text generations are aligned with the reward model learned over preferences:
@@
For the purpose of this paper we only focus on single-turn dialogue where only a single action is taken by the model so only consider the immediate reward as this is the most common RLHF setting we have datasets for. Future work should consider multi-step RL settings such as those found in Pan et al. (2023); Andriushchenko et al. (2024) where we’d need to consider a discounted cumulative reward and provide an analysis of the underlying Markov Decision Process (MDP).
A Reverse Preference Attack (RPA) is performed by reversing (“flipping”) the preference ordering—for example, by switching labels of preferred and dis-preferred over text samples in a binary preference dataset—inducing a ”reverse” reward model $R_{\phi}^{\text{reverse}}(y,x)$. If the preference data is primarily over harmlessness, then this reward model can be used as an objective for learning harmful text generation policies by downstream RLHF algorithms. While we primarily provide experiments where all of these preferences have been reversed in order to understand the worst-case capabilities of current defences, in practice it is more likely that an adversary attacking an online feedback collection mechanism or dataset may only switch a small number of labels to avoid getting caught. This threat model is similar to Rando and Tramèr (2023) except that we don’t focus on an implanted backdoor but on preference order reversal. Yi et al. (2024) introduced a special case of RPA in a restricted DPO setting. They showed “reverse DPO” was effective in undoing safety guards of LLMs. However, they neglected providing a full framework accounting for the vulnerability of LLMs under adversarial feedback.
Why should fine-tuning or preference learning not compromise safety and why are HFTAs and RPAs consider attacks? While we should expect that current LLMs learn to model the preferences of any dataset given to them, we believe there is argument to be made that this is a vulnerability. Echoing (Rosati et al. 2024b), if safety guards can easily be removed through fine-tuning (or in this case preference learning) then can we call these models safe? Should we consider open-weight release responsible given these known vulnerabilities? Finally, we are hopeful that the empirical evidence from Tamirisa et al. (2024); Rosati et al. (2024b); Zou et al. (2024) that defences is possible.
Our results and analysis shows that RPAs are a much more limited threat than SFT due to the complexities of RLHF, and as we will show later, that RPAs are a weaker attack. However the purpose of our work is not to introduce a new stronger attack but to analyzing the vulnerability of LLMs under feedback collection. Additionally, it may be the case that the only means of attack is through providing feedback or a feedback dataset. Regardless of the significance of this threat, we believe RPAs are an important tool to highlight the vulnerability of LLMs during RLHF especially given the risks elucidated in Casper et al. (2023) which highlights that unsafe feedback may be collected by accident, due to annotator misunderstanding, or may be a natural outcome of disagreements about what alignment means across communities. Finally, we believe that this work could provide early insights on a much more realistic vulnerability, LLM agents learning online in a full RL setting such as what has been presented in Pan et al. (2023); Andriushchenko et al. (2024).
Vulnerability Analysis
@@
To illustrate the vulnerability of LLMs under RPAs, we assume a setting where the attacker has the capability to perform RLHF on an aligned language model using their own data—either by acquiring the weights of a model, using a preference learning API, or feedback collection mechanism exposed through by an LLM provider. We perform our experiments on the safety-aligned open-weight model llama2-7b-chat (Touvron et al. 2023) since there are no preference learning APIs exposed for closed models and there is no way to perform a controlled experiment using their feedback collection mechanisms. As attack methods, we use Harmful Fine-tuning Attacks (Rosati et al. 2024b) (using supervised fine-tuning (SFT) on a harmful dataset), and RPAs (using DPO and PPO). As attack datasets, we use the harmlessness preference datasets of BeaverTails (Ji et al. 2023) and Safe RLHF (Dai et al. 2024). We flip a certain percentage of binary harmfulness preference labels of 1,000 randomly picked samples, and use each training algorithm to further train the original aligned LLM. A varying ratio of preference labels are flipped in order to simulate either attacker stealthiness or label noise. This experiment is similar in spirit to Rando and Tramèr (2023) except that we are not introducting a backdoor. $1,000$ samples is small considering how large many RLHF datasets are and can be used for training models on a limited budget.
@@
Table 1 presents experimental results for the case when 100% of the labels in an attack dataset are flipped. Specifically, it shows the mean probability of the harmfulness label of the answer to a held-out evaluation set of 100 harmful questions from each attack dataset using a harmfulness classifier. We observe that all three training algorithms are able to break safety alignment in the original model to a considerable degree on as little as 1,000 attack samples—RPA using DPO being the most effective. Note that since we use the likelihood of assigning a harmfulness label, a mean score of over 0.5 indicates that a safety guard is effectively removed as a majority of answers are now classified as harmful. DPO and PPO use policy divergence constraints, $\beta=0.1$ and $kl=0.2$ respectively, with the original policy, i.e. the aligned model. We vary these constraints in the Appendix (Analysis of Reference Model Divergence) and find that policy divergence constraints with an aligned model are not an effective defence for DPO but may be effective for PPO which is likely because KL is used as a more direct constraint for PPO. Details of the experimental conditions of the attack, harmfulness measure and classifier, and datasets used are presented in the Appendix (Implementation Details).
Dataset No Attack SFT DPO PPO
BeaverTails 0.06 0.68 0.70 0.69
Safe RLHF 0.07 0.75 0.83 0.59
@@
Next (Figure 2 below), we present the setting where labels are flipped at a given ratio $\rho\in[0,0.9]$ to simulate two settings: (1) unintended data corruption due to label noise, and (2) adversarial data poisoning where an attacker hides a given number of labels. In the case of SFT, label flipping means we are using a harmful sample for causal language modelling instead of a harmless one. We use the same experimental setup as above, except that we also employ an unaligned model llama2-7b in order to understand the effects of vulnerability to HFTAs and RPAs during during safety training as these datasets are commonly used for that purpose. As shown in Table 2, in as little as 25% label flips for Safe RLHF we have successful attacks (DPO - 0.51) which indicates that these models may be vulnerable to label noise or stealthy attacks. However, we acknowledge in practice a corruption of 25% of the dataset is likely detectable and easily mitigated through filtering by a model developer–we explained earlier why we should still take this vulnerability seriously. RPAs using DPO remains the most effective attack across noise levels and both datasets. Observe that DPO’s and PPO’s harmfulness does not always increase linearly with attack strength which due to the instability of training. As expected unaligned models are even more vulnerable than aligned ones which indicates that noisy harmlessness datasets could introduce significant issues during alignment training and that while not generally effective as a mitigation strategy, pre-existing safety guards can make attacks more difficult to perform.
@@
Figure 2: Vulnerability of models under varying ratios (the columns) of harmless preference flips. The mean harmfulness scores are presented as above. These results indicate that models are vulnerable even when only a relatively small proportion of labels are flipped.
Defences as Learning Constraints
To introduce defences against RPAs, we introduce Constrained Markov Decision Processes (Altman 2021) which have been actively studied for modelling Safe Reinforcement Learning (Gu et al. 2024a) and recently in the RLHF for LLM setting (Dai et al. 2024). Formally, an CMDP is the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\mathcal{C})$ [^1] Since we are focused on single turn dialogue, we don’t present the typical starting state distribution and discount factor. consisting of a set of states $\mathcal{S}$, a set of actions $\mathcal{A}$, a transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ that denotes the probability $\Pr(s^{\prime}|s,a)$ of entering the next state $s^{\prime}\in\mathcal{S}$ given the agent is in a previous states $s\in\mathcal{S}$ and takes action $a\in\mathcal{A}$. Like MDPs, we also have a reward function $\mathcal{R}:\mathcal{S}\times{A}\rightarrow\mathbb{R}$ which provides a scalar reward after taking an action $a$ when in state $s$. Unique to CMDPs is the constraint set $\mathcal{C}=\{c_{i},b_{i}\}_{i}^{N}$ which is a set of cost functions $c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ that we require are above the threshold $b$.
@@
Key to our analysis is that the set of policies $\Pi_{\Theta}$ that are learnable in a typical policy learning algorithm, i.e. Eq 1, are restricted by a cost function $c$ and a cost threshold $b$ where the acceptable policy set is in $\Pi_{c}=\{\pi_{\theta}\in\Pi_{\Theta}|\>c(s,a)\leq b,\forall(s,a)\sim\pi_{\theta}\}$. When the cost function is a harmfulness measure and the threshold is set to allow a degree of harm that is acceptable, this definition adapts the resistance condition for defences against HFTAs from Rosati et al. (2024b) to the RL setting. We can connect this further to the attacker’s budget by using a measure of how long $\pi_{\theta}$ remains in $\Pi_{c}$ across training steps $t$ where the defender wants to prolong policy constraints under an attack.
A viable defence against a RPA must solve a learning problem where constraints of the CMDP are satisfied while optimizing a given reward function. For the scope of this paper, we focus our analysis on empirical observations of whether proposed defences ensure the LLM policy stays in the acceptable policy set $\Pi_{c}$ by measuring the cost function $c$ and ensuring it is a low as possible. Future work should provide theoretical analysis of whether defences solutions solve CMDPs optimally. Finally, we acknowledge the existence of many algorithms in safe RL broadly that have both well establish theoretical and empirical results that are reviewed in Gu et al. (2024b); Brunke et al. (2022); Garcıa and Fernández (2015) but are beyond the scope of this paper since (1) adaptation to an LLM settings requires non trivial modifications and (2) these methods are not directly applicable as defences against harmful fine-tuning attacks which we want to illustrate as a control experiment.
@@
Instead, we evaluate previously proposed defences against harmful fine-tuning attacks and jailbreaking in the LLM context. The only exception is the Safe RLHF method (Dai et al. 2024) which uses a similar approach to previous attempts at solving CMDPs with primal-dual methods. In Table 2 we classify defence candidates into two categories by defender capability assumptions: online and offline. Online defences assume that the defender does not have control of the training data but does have control over the training process. Online defences work by adopting different variants of explicit constraints on the training process. On the other hand, offline defences assume that the defender has applied before publicly releasing the weights of an LLM and can make no other interventions. Offline defences use implicit constraints [^2] Terminology borrowed from Boyd and Vandenberghe (2004, p.134) that exist in the model weights themselves as demonstrated by their ability to prevent defended models from learning policies outside of the constraint set. Each defence is described in detail in the Appendix (Defences)  [^3] Other viable defence settings such as data filtration or post-training output monitoring are viable but outside of our scope of interest.
Constraint Defence Paper
Embeddings Vaccine Huang, Hu, and Liu2024
Weights Lisa Huang et al.2024
Rewards Safe RLHF Dai et al.2024
Loss Security Vectors Zhou et al.2023
Loss Refusal Loss Qi et al.2024a
Representation RepNoise Rosati et al.2024a
Representation Circuit Breakers Zou et al.2024
Representation RMU Li et al.2024b
Meta-learned TAR Tamirisa et al.2024
@@
Table 2: Taxonomy of defences candidates against RPAs showing that defences can be analyzed based on the learning constraints they provide. The constraint column indicates how the harmlessness constraint is implemented. Defences above the line “online” or “explicit” defences while below the line are “offline” or “implicit” methods.
@@
Table 3: Evaluation of online defences to HFTA and RPA variants. Missing metrics, indicated by dashes, indicate the training process could not complete due to sampling errors. Refusal loss and Lisa maintain the lowest harmfulness and highest helpfulness. Asterix indicates a disfluent model which can be seen in the PPL scores.
Online Defences
We evaluate the online defences in Table 3 on the same attacks in the above vulnerability analysis. In addition to mean harmfulness scores, we use three additional metrics:
  1. 1.
    Helpfulness, measured by taking 100 samples of the preferred helpful questions from the Anthropic-HH helpfulness dataset (Bai et al. 2022) and using a helpfulness reward model (details in the Appendix: Implementation Details) to rate generated responses to these questions. This measure is introduced to understand the degree to which the defence maintains general capabilities on harmless tasks.
  2. 2.
    Perplexity (PPL), measuring the fluency of the helpfulness answer above using gpt-2 (Radford et al. 2019).
  3. 3.
In Table 3 we find that across all methods Lisa and Refusal Loss are the most effective defences, maintaining low perplexity and high helpfulness. Security Vectors and Vaccine provide some defence against HFTA using SFT, but they do so at the cost of lower helpfulness scores. They also do not work in the other settings: Security Vectors is successfully attacked under RPA using both PPO and DPO, and Vaccine results in sampling errors during RPA/PPO due to numeric instability of the learned log probabilities or results in complete degeneration in the case of DPO, as evidenced by the method’s perplexity. For Safe RLHF, the method is limited to the PPO setting only as it is a reward shaping method and is not as effective (Safe RLHF dataset) or successfully attacked (BeaverTails). Interestingly, successful defences often have a very high KL divergence with the original policy. While KL divergence magnitudes are not necessarily informative of the actual distributional distances, since it is not a distance metric, our finding indicates that KL divergence with respect to the original policy model is not an effective indicator of preventing the undoing of safety alignment.